# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

### v3

@@ 5 Ablation
Table 1: Instruction-Following Results. For each model (columns), we report the pre-trained and FIT average GPT-40 ratings, and the two-sided Wilcoxon Signed-Rank $p$-value testing the difference between the distributions of ratings.
To test for significant differences in instruction-following performance, we conducted a two-sided Wilcoxon Signed-Rank Test (Wilcoxon, 1992) on paired ratings. The null hypothesis assumes no difference in the median ratings between pre-trained and FIT models. The results, summarized in Table 1, show no statistically significant differences ($p>0.05$) in instruction-following performance across all models. These findings demonstrate that FIT maintains general instruction-following capabilities whilst additionally enhancing test-time steerability.
@@ 6 Conclusion
In this work, we introduce Focus Instruction Tuning (FIT), a method designed to steer the behaviour of LLMs by focusing on or ignoring specific features when formulating responses. Across a range of tasks and settings, we demonstrate that FIT provides dynamic and precise control over LLM behaviour at inference time, enabling users to adapt model responses even in the context of distribution shifts over feature values or when generalising to unseen features. Furthermore, our approach can address challenges such as mitigating the influence of known stereotypes that might otherwise impact responses, showcasing one of its many applications. Thus, FIT represents a step toward enabling more robust, steerable, fair, and controllable LLMs.
Appendix A Impact Statement
@@ Appendix B Limitations and Future Work
  • Regulatory and Ethical Expectations: Regulatory frameworks and ethical guidelines increasingly require the explicit identification and mitigation of problematic features (of the European Parliament, 2016) Corresponding initiatives aim to define and enforce measurable categories of “violating behaviour” in AI models. By providing a mechanism to steer model behaviour based on these identified features, FIT effectively complements efforts to promote fair and transparent predictions (Guldimann et al., 2024; Zeng et al., 2024).
@@
  • Compatibility with Automated Spurious Feature Identification: FIT also works seamlessly with automated methods for detecting spurious features (Wang & Culotta, 2020; Wang et al., 2022; Zhou et al., 2023; Zheng et al., 2024). Whether spurious features are labelled manually or derived from algorithmic detection, they can be harnessed by FIT’s focus instructions at inference time. This compatibility enables a comprehensive approach to managing known issues and responding to newly uncovered features as they arise.
@@
Scope of Experiments and Extensions to Open-Ended Tasks. Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Extending FIT to open-ended tasks, such as summarisation or translation, remains an exciting direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Overlapping Features and Ambiguities. Additionally, our evaluation on the HANS dataset Appendix L revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
Appendix C FT Training and Optimisation Settings
FT Optimisation. Algorithm 1 gives precise details on how we implement FIT in practice when performing ERM of a model on a given training set. In particular, it shows how we approach optimising the FIT training objective given in Equation 8.
@@ Appendix C FT Training and Optimisation Settings
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, The feature set contains $\mathcal{F}$, instruction $I$, model parameters $\theta$, batch size $B$, number of epochs $E$, step size $\eta$, and mapping function $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},y,s)$.
2:  Initialise: Model parameters $\theta$, optimiser
@@
7:           Sample focus instruction $I_{\text{focus}}^{b}\sim p_{\mathcal{I}_{\text{focus}}}$,
@@
$\ell(\theta)=\frac{1}{B}\sum_{b=1}^{B}-\log p_{\theta}(y_{\text{focus}}^{b}|I,I_{\text{focus}}^{b},x^{b})$
@@
$\theta\leftarrow\theta-\eta\nabla_{\theta}\ell(\theta)$
@@
FT training settings. We use the SFTTrainer class from HuggingFace (Wolf et al., 2020) and use all of the default training settings for performing SFT of LLMs. Furthermore, we define $p(\mathcal{I}_{\text{focus}})$ by placing a small probability (in our experiments, 0.05) on the empty focus instruction $\emptyset$. We then uniformly distribute the remaining probability mass over the non-empty focus instructions.
We implement early stopping on a held-out validation set based on the cross-entropy loss over focus labels $y_{\text{focus}}$ corresponding to randomly sampled focus instructions - this matches the context in which the models will be evaluated. We obtain this set by splitting our training set in a 90/10% ratio for training and validation splits respectively. We use a patience of $4$ validation evaluation steps, which occur after a fixed number of steps.
@@
We use LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. We target the query and value projection matrices within each LLM and use LoRA $r=16$ and $\alpha=32$ across models.
@@
Choice of $\rho_{\text{spurious}}$ during training. In our synthetic experiments, we set up a controlled environment by imposing two independence conditions: $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$. These ensure that (i) the ground-truth label cannot be predicted using the spurious feature $S$, and (ii) the spurious label cannot be predicted using the core feature $C$. By removing direct correlations between these features and labels, the model is leans to focus on the specified feature, without being influenced by the other feature, avoiding any potential shortcuts that could be exploited if these conditions did not hold.
@@
  • Independence $Y\perp\!\!\!\perp S$: This condition prevents the model from leveraging spurious feature $S$ to predict ground-truth label $Y$. With no predictive signal from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario enabling prediction only through the specified feature indicated through the focus instruction passed to the model.
@@
  • Independence $Y_{S}\perp\!\!\!\perp C$: This condition serves a complementary role in preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions alone for making predictions.
@@
While these conditions represent an ideal setting, they are not strictly necessary for FIT to work in practice. Indeed, real-world data rarely satisfies such perfect independence, and we illustrate the robustness of the method in more realistic scenarios through our BBQ experiments in Section 4.3 where correlations between $Y$ and $S$ or between $C$ and $Y_{S}$ may exist as no subsampling or dataset manipulations have been made. By examining both the controlled environment and more naturalistic datasets, we demonstrate that our approach can handle scenarios with varying degrees of spurious correlations.
@@
To achieve this independence in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Appendix I and Appendix K, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
Appendix D Evaluation Metrics
@@
Generation settings. We generate responses from our FT model using constrained beam-decoding (Anderson et al., 2017) with $8$ beams. This ensures that the answer labels for each classification task that we investigate appear in the model’s output. We limit the maximum number of newly generated tokens to be $5$ to stop any unnecessary text given after the model’s initial classification prediction.
@@
Computing the focus accuracy metric. We report the focus accuracy $\mathcal{A}_{\text{focus}}$ of generations when evaluating FT models. As we are guaranteed to include the task labels within the model’s response through constrained decoding, we simply check to see if the focus label, $y_{\text{focus}}$, is within the model’s response or not in order to determine if the model’s response is correct.
Appendix E FIT Focus Instructions and Prompt Templates
Prompt templates. Here we provide the prompt templates that we use for FIT. Note, that we drop the feature considerations section and focus instruction from the prompt to form the instruction prompt $\emptyset$ with no focus instruction.
SS Focus instruction prompt templates $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model performing sentiment analysis on a binary dataset, making predictions from the labels [negative, positive]. Make your prediction based on the relevant features described below, focusing on the specified features and ignoring those deemed irrelevant. For the input below, output either negative or positive ONLY for your prediction of the input’s label.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### Input: {input}
### Output:
NLI focus instruction prompt template $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model performing natural language inference (NLI) on a dataset with three labels [entailment, neutral, contradiction]. For the inputs consisting of premise and hypothesis below, make your prediction based on the relevant features described, focusing on the specified features and ignoring those deemed irrelevant. Output either entailment, neutral, or contradiction ONLY for your prediction of the input’s label.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### Premise: {premise}
### Hypothesis: {hypothesis}
### Output:
BBQ focus instruction prompt template $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model answering a QUESTION based on a provided CONTEXT, choosing to ANSWER with one of three options: 1, 2, or 3. Each option corresponds to an answer from the provided ANSWER OPTIONS. You must output EXACTLY one of these numbers (1, 2, or 3) representing the answer you select.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### CONTEXT: {context}
### QUESTION: {question}
### ANSWER OPTIONS:
(1) {ans0}
(2) {ans1} (3) {ans2} ### ANSWER:
Focus instructions. We consider the following focus instruction formats for the different focus instructions introduced in Equation 2 which are used for FIT training and evaluation:
Focus instructions $\mathcal{I}_{\text{focus}}$ For features $F_{i},F_{j}\in\mathcal{F}$:
Focus instructions $\text{focus}(F_{i})$:
Direct your attention solely to $F_{i}$. Concentrate all your reasoning on $F_{i}$. Make $F_{i}$ the central factor in your decision. Base your judgment exclusively on $F_{i}$. Pay attention only to $F_{i}$ when making your prediction. Use $F_{i}$ as the key input for your evaluation. Focus entirely on $F_{i}$ and ignore other aspects. Rely exclusively on $F_{i}$ to reach your conclusion. Consider only $F_{i}$ and disregard all else. Let $F_{i}$ be the primary basis for your decision.
Ignore instructions $\text{ignore}(F_{i})$: Completely rule out $F_{i}$ from your reasoning. Disregard any influence of $F_{i}$ in your prediction. Treat $F_{i}$ as irrelevant to your decision-making process. Exclude $F_{i}$ entirely from your evaluation. Do not let $F_{i}$ play any role in your assessment. Intervene to prevent $F_{i}$ from affecting your prediction. Ensure that $F_{i}$ has no bearing on your final decision. Block $F_{i}$ from contributing to your reasoning. Negate the impact of $F_{i}$ in your prediction. Ruling out $F_{i}$ is crucial—do not let it affect your decision.
Focus and Ignore instructions $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ Focus specifically on $F_{i}$. Disregard $F_{j}$ in your decision-making process. Base your prediction solely on $F_{i}$. Exclude $F_{j}$. Direct all your attention to $F_{i}$. Block out $F_{j}$ from your prediction. Consider only $F_{i}$ in your reasoning. Rule out $F_{j}$ in your decision-making. Prioritize $F_{i}$. Completely ignore $F_{j}$ in your prediction. Do not consider $F_{j}$ in your decision-making process. Focus exclusively on $F_{i}$. Ignore any influence of $F_{j}$. Concentrate on $F_{i}$ in your prediction. Disregard $F_{j}$ entirely. Base your analysis solely on $F_{i}$. Rule out $F_{j}$ in your prediction. Shift your focus to $F_{i}$. Do not pay attention to $F_{j}$ in your decision-making process. Rely only on $F_{i}$.
Appendix F Additional Baselines Results

### v4

@@ 5 Ablation Studies
Table 1: Instruction Following After FIT. For each base model (columns), we report the pre-trained and FIT average GPT-40 ratings, and the two-sided Wilcoxon Signed-Rank $p$-value testing the difference between the distributions of ratings.
Zero-Shot Transfer (MMLU).
We next measure zero-shot transfer to MMLU (Hendrycks et al., 2021), a MCQA dataset, testing problem solving and general world knowledge of models. Table 2 shows accuracy and perplexity (across entire model vocabulary) for pre-trained and FIT Llama and Mistral models. FIT changes accuracy by at most 0.8%, while showing lowering perplexity, indicating that task performance of pre-trained models has not been sacrificed through FIT. This demonstrates that FIT does not hurt existing transfer performance of base models in a zero-shot setting.
Model Llama Mistral
Pre-Trained FIT Pre-Trained FIT
Accuracy  ($\uparrow$) 30.4 29.6 29.4 29.0
Perplexity ($\downarrow$) 6.29 2.79 15.2 5.22
Table 2: Zero-Shot MMLU After FIT. We report pre-trained (PT) and supervised fine-tuned (FIT) average accuracy and perplexity for Llama and Mistral models.
5.4 Model Size Ablation
We further assess FIT across the Qwen-2.5-Instruct (Yang et al., 2024a) family of models on three scales (1.5B, 3B and 7B parameters) on the BBQ dataset under the same training conditions as in Section 4.2. As shown in Figure 8, FIT shows strong steerability across all model sizes, even for the smallest 1.5B model. Moreover, we see that performance generally scales with model size. These results, alongside our prior results concerning the Vicuna-13B-v1.5 model demonstrate FIT’s robustness to model capacity and its favourable scaling behaviour.
Figure 8: Model Size Ablation. Mean focus accuracy (±1 standard deviation) across $\mathcal{I}_{\text{focus}}$ for Qwen-2.5-Instruct models at 1.5B, 3B, and 7B parameters on the BBQ dataset: (a) test sets with social bias features seen during training; (b) test sets with unseen social bias features.
@@ 6 Conclusion
In this work, we introduce Focus Instruction Tuning (FIT), a method designed to steer the behaviour of LLMs by focusing on or ignoring specific features when formulating responses. Across a range of tasks and settings, we demonstrate that FIT provides dynamic and precise control over LLM behaviour at inference time, enabling users to adapt model responses base on natural language instructions. This steering behavoi holds even in the context of distribution shifts over feature values or when generalising to unseen features. Furthermore, our approach can address challenges such as mitigating the influence of known stereotypes that might otherwise impact responses, showcasing one of its many applications. Thus, FIT represents a step toward enabling more robust, steerable, fair, and controllable LLMs.
Acknowledgments
This work is supported by the UKRI grant: Turing AI Fellowship EP/W002981/1. Alasdair Paren is supported by UK ASIS Systemic AI Safety Grant. Francesco Pinto acknowledges support through the DSI, University of Chicago. Adam Davies is supported by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education, through Award #2229612 (National AI Institute for Inclusive Intelligent Technologies for Education). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of National Science Foundation or the U.S. Department of Education.
Impact Statement
@@ Acknowledgments
Appendix A Author Contributions
  • Tom A. Lamb:
    • $\circ$
      Idea: Turned the initial project ideas and directions into the concrete methodology described in Section 3.
    • $\circ$
      Experiments and Methods: Designed and carried out all experiments in the paper. These are described in Figure 3, Figure 5, Appendix E and Section 5.
    • $\circ$
      Dataset Creation: Designed and carried out the synthesis of the SS and SMNLI dataset presented in Appendix E and Section 4.1.
    • $\circ$
      Analysis: Led the analysis of all empirical results.
    • $\circ$
      Paper Writing: Led writing of all sections (main paper and appendix), including both the initial drafts of the paper and the final version of the paper presented here; and led literature review on spurious feature learning, instruction-tuning and helped with the aligning LLMs section of Section 2.
  • Adam Davies:
    • $\circ$
      Idea: Co-conceived the research problem and core approach of FIT (with Francesco Pinto).
    • $\circ$
      Supervision: Led advising of methodology and experimental design.
    • $\circ$
      Analysis: Advised and assisted with analysis of all empirical results.
    • $\circ$
      Paper Writing: Co-wrote drafts of Abstract, Section 1 and Section 2; led literature review for mechanistic interpretability, aligning LLMs, and latent steering in Section 2, and contributed to literature review for spurious feature learning and instruction tuning; and helped edit and revise all sections of the paper.
  • Alasdair Paren:
    • $\circ$
      Supervision: Co-advised methodology and experimental design with a particular focus on the simplicity bias and shortcut learning.
    • $\circ$
      Paper Writing: Provided feedback on drafts.
  • Philip Torr:
    • $\circ$
      Supervision: Provided general advice on initial project direction; helped secure compute resources for the experiments conducted in this paper.
    • $\circ$
      Paper Writing: Provided draft feedback on the initial sections of the paper.
  • Francesco Pinto:
    • $\circ$
      Idea: Co-conceived the research problem and core approach of FIT (with Adam Davies).
    • $\circ$
      Supervision: Co-advised methodology and experimental design.
    • $\circ$
      Paper Writing: Provided feedback on earlier versions of the drafts of the paper.
@@ Appendix B Limitations and Future Work
  • Regulatory and Ethical Expectations: Regulatory frameworks and ethical guidelines increasingly require the explicit identification and mitigation of problematic features (of the European Parliament, 2016). Corresponding initiatives aim to define and enforce measurable categories of “violating behaviour” in AI models. By providing a mechanism to steer model behaviour based on these identified features, FIT effectively complements efforts to promote fair and transparent predictions (Guldimann et al., 2024; Zeng et al., 2024).
@@
  • Compatibility with Automated Spurious Feature Identification: FIT also works seamlessly with automated methods for detecting spurious features (Wang & Culotta, 2020; Wang et al., 2022; Zhou et al., 2024b; Zheng et al., 2024). Whether spurious features are labelled manually or derived from automated detection, they can be harnessed by FIT’s focus instructions at inference time. This compatibility enables a comprehensive approach to managing known issues and responding to newly uncovered features as they arise.
@@
Scope of Experiments and Extensions to Open-Ended Tasks. Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Whilst we provide positive evidence through our BBQ-NLG ablation in Section 5.1, extending FIT to open-ended tasks such as summarisation or translation, remains an key direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Overlapping Features and Ambiguities. Additionally, our evaluation on the HANS dataset Appendix K revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
Handling Multiple Features. While our current experiments provide initial evidence that FIT can handle combined instructions, specifying both a feature to focus on and a feature to ignore, we recognise the importance of scaling this capability. An important direction for future research is extending FIT to manage instructions involving multiple features simultaneously, including cases where models are required to focus on or ignore several attributes in tandem.
Appendix C FIT Focus Instructions and Prompt Templates
Prompt Templates. In Figure 10 we provide the prompt templates that we use for FIT training and for general evaluation across all methods. Note, that we drop the feature considerations section and focus instruction from the prompt to form the default prompt $\emptyset$ introduced in Equation 2.
Focus Instructions. In Figure 9 we show the focus instruction formats for the different focus instructions introduced in Equation 2, defining $\mathcal{I}_{\text{focus}}$ throughout our experiments.
Figure 9: Focus Instructions. Focus instructions that are used for focussing and ignoring features $F_{i},F_{j}\in\mathcal{F}$ during FIT training and for general evaluation.
Figure 10: FIT Prompt Templates. Prompt templates used for FIT training and evaluation across all four datasets that we investigate over.
Appendix D FIT Training, Optimisation and Evaluation
D.1 FIT Training and Optimisation
FT Optimisation. Algorithm 1 gives precise details on how we implement FIT in practice when performing ERM of a model using the FIT training objective given in Equation 3 on a given training set.
@@ D.1 FIT Training and Optimisation
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, The feature set $\mathcal{F}$, focus instructions $\mathcal{I}_{\text{focus}}$, instruction $I$, model parameters $\theta$, batch size $B$, number of epochs $E$, step size $\eta$, and focus label mapping $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},y,s)$.
2:  Initialise: Model parameters $\theta$, optimiser.
@@
7:           Sample focus instruction $I_{\text{focus}}^{b}\sim p_{\mathcal{I}_{\text{focus}}}$.
@@
$\ell(\theta)=\frac{1}{B}\sum_{b=1}^{B}-\log p_{\theta}(y_{\text{focus}}^{b}|I,I_{\text{focus}}^{b},x^{b}).$
@@
$\theta\leftarrow\theta-\eta\nabla_{\theta}\ell(\theta).$
@@
FT Training Settings. We use LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. We target the query and value projection matrices within each LLM and use LoRA $r=16$ and $\alpha=32$ across models.
We implement early stopping on a held-out validation set based on the cross-entropy loss over focus labels $y_{\text{focus}}$ corresponding to randomly sampled focus instructions - this matches the context in which the models will be evaluated. We obtain this set by splitting our training sets in a 90/10% ratio for training and validation splits respectively. We use a patience of $4$ validation evaluation steps, which occur after a fixed number of steps.
During training, we define $p(\mathcal{I}_{\text{focus}})$ by placing a small probability (in our experiments, 0.05) on the empty focus instruction $\emptyset$. We then uniformly distribute the remaining probability mass over the non-empty focus instructions.
@@
Choice of $\rho_{\text{spurious}}$ During Training. In our synthetic experiments described in Figure 3 and Appendix E, we set up a controlled environment by imposing two independence conditions: $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$. These ensure that (i) the ground-truth label cannot be predicted using the spurious feature $S$, and (ii) the spurious label cannot be predicted using the core feature $C$. By removing direct correlations between these features and labels, the model learns to focus on the specified feature, without being influenced by another feature, avoiding any potential shortcuts that could be exploited if these conditions did not hold.
@@
  • Independence $Y\perp\!\!\!\perp S$: This condition prevents the model from leveraging spurious feature $S$ to predict ground-truth label $Y$. With no predictive signal from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario. This enforces prediction only through the specified feature indicated through the focus instruction passed to the model.
@@
  • Independence $Y_{S}\perp\!\!\!\perp C$: This condition serves a complementary role in preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions for making predictions.
@@
While these conditions represent an ideal setting, they are not strictly necessary for FIT to work in practice. Indeed, real-world data rarely satisfies such independence assumptions. We illustrate the robustness of the method in more realistic scenarios through our BBQ experiments in Section 4.2 where correlations between $Y$ and $S$ or between $C$ and $Y_{S}$ may exist, as no subsampling or dataset manipulations have been made to enforce these conditions. By examining both controlled environments and more natural datasets, we demonstrate the robustness of FIT.
@@
To achieve the aforementioned independence assumptions in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Appendix E and Appendix F, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
D.2 Evaluation
@@
Generation Settings. We generate responses from models using constrained beam-decoding (Anderson et al., 2017) with $4$ beams. This ensures that the answer labels for each classification task that we investigate appear in the model’s output. We limit the maximum number of newly generated tokens to be $5$ to stop any unnecessary text given after the model’s initial classification prediction.
@@
Computing the Focus Accuracy Metric. We report the focus accuracy $\mathcal{A}_{\text{focus}}$ of generations when evaluating models. As we are guaranteed to include the task labels within the model’s response through constrained decoding, and have reduced the maximum number of tokens that a model can generate at inference-time, we simply check to see if the focus label, $y_{\text{focus}}$, is within the model’s response or not in order to determine if the model’s response is correct or not.
D.3 Dataset Sizes
The sizes of each of the splits of the SS, SMNLI and BBQ datasets are given in Table 3.
Table 3: Dataset Sizes. Dataset split sizes for SS, SMNLI and BBQ datasets.
SS SMNLI BBQ
Training 5296 7200 16700
Validation 1324 1800 1590
Test 1818 900 2352
D.4 Complete Set of Baselines